Find the power function that the graph of f resembles for large values of |x|. f(x)=(x+5)^2
Answer by jim_thompson5910(35256)

f(x)=(x+5)^2

f(x) = (x+5)*(x+5)

f(x) = x^2 + 5x + 5x + 25 ... use FOIL

f(x) = x^2 + 10x + 25

The leading term is x^2

The final answer is g(x) = x^2 (note: your teacher may want you to drop the g(x) notation. It will depend on the teacher so it's best to ask)